Skip to Main Content
PTAGIS logo link back to home page

Interrogation Site Metadata

BCC - BON PH2 Corner Collector

Status:   Active - Operational
Site Type: Juvenile Fish Bypass Facility  |  Organization: Pacific States Marine Fisheries Commission
General Operational Period: Seasonally
Start Date: 4/13/2006 8:00 AM

This site monitors the Bonneville Dam second powerhouse corner collector outfall channel. It consists of one large antenna through which the channel flows. This site is operated seasonally when the corner collector is used to pass large debris.

Current Operations

Last File Loaded: BCC-2022-176-P-019.xml
Date Last File Loaded: 6/25/2022 9:50 PM
Date Last File Opened: 6/25/2022 6:56 PM
Date Last File Closed: 6/25/2022 7:56 PM
Last Observation Tag:
Last Observation Date:
Detection Summary: Last Day's Observations

Contacts

Primary Contact:

Don Warf
dlwarf@psmfc.org
509-735-2773

Data Steward:

Nicole Tancreto
ntancreto@psmfc.org
503-595-3141

Location

Coordinates: 45.648140, -121.942544  |  River Km: 234

Operational History

Operational history describes the period the site was interrogating for PIT tags. The comments provide anecdotal information about the operation of the site during this period. Site operational periods are updated typically at the end of the calendar year.

Year Start Date End Date Comments
2022
2021 3/26/2021 12:00 AM 9/1/2021 12:00 AM
2020 3/20/2020 12:45 PM 9/1/2020 12:00 AM
2019 4/9/2019 5:05 AM 9/1/2019 12:00 AM
2018 4/7/2018 1:37 AM 8/31/2018 11:00 PM
2017 3/30/2017 12:00 AM 9/1/2017 12:00 AM
2016 3/5/2016 12:00 PM 8/31/2016 12:00 AM
2015 2/13/2015 12:00 AM 9/1/2015 12:00 AM Channel was dewatered February 24-March 10. Channel was dewatered for the year on September 1.
2014 3/17/2014 12:00 AM 12/15/2014 12:00 AM Site was dewatered August 31-November 13
2013 3/1/2013 12:00 AM 10/31/2013 2:05 PM Channel was watered down between September 1 and October 11. It was operated intermittently from October 11-31 primarily at night to facilitate USACE repairs and PIT tag anntenna testing.
2012 3/19/2012 7:18 AM 9/16/2012 12:11 AM The channel watered up on March 19 and was dewatered on September 16.
2011 3/3/2011 7:20 AM 9/1/2011 7:30 AM The B2CC was open from 0731 until 1547 on 09 February to flush debris; the PIT tag detector was not active during this brief interval. B2CC watered up for two days, March 3-4, for kelt fallback testing; the BCC transceiver was active during these tests. The B2CC flume was then dewatered from March 4 until March 14. The BCC channel was dewatered for the season on September 1.
2010 3/14/2010 6:45 AM 8/31/2010 3:00 PM The corner collector flume watered-up for the 2010 season on March 14. It was shut down for approximately 20 hours from 10:40 on March 16 until 7:00 on March 17. The flume was dewatered for 54 hours from 6AM PST on April 20 until noon (PST) on April 22. The flume was closed for 20 hours from April 26 at 12:30 PM (PST) to April 27 at 9:45 AM due to a broken BGS cable and no boat protection. The flume was dewatered at the end of the spill season on August 31.
2009 4/3/2009 6:30 AM 9/3/2009 8:00 AM The BCC flume watered up on April 3 and operated without operation until Sept. 1, when dewatering activities began. The flume was fully dewatered on Sept. 3.
2008 3/6/2008 12:00 PM 9/1/2008 6:50 AM The Bonneville PH2 Corner Collector channel (B2CC) watered-up briefly a few times between late February and early March to pass debris, but was not operated consistently, nor was the transceiver initialized, until March 6. The Corps regularly dewatered the B2CC between March 6-21 to allow researchers to get to their equipment. The B2CC was dewatered, and the BCC transceiver disabled, between March 21 at 13:35 and March 24 at 9:04. The B2CC was dewatered at the end of the fish passage season on September 1.
2007 3/1/2007 1:30 PM 8/31/2007 1:00 PM The Corner Collector channel watered up on March 1, 2007. The channel was dewatered for the year on August 31.
2006 4/13/2006 8:00 AM 11/21/2006 9:00 AM New installation. First year. Opened for the season at 0815 on 13 April. The channel was shut down for the season around 16:00 PDT on 8/31/06. However, PIT tags were somehow detected through noon on September 1. The bypass channel was unexpectedly re-watered at 10:45AM on November 15, to pass trash and debris in the BON-PH2 forebay. The PIT tag system was re-activated at 11:30am on Nov. 16. The bypass was dewatered at ~8:30am on Nov. 21. The PIT tag detector was active during this interval, but no tags were detected. The B2CC was watered up between 12/21-26/06 to pass debris during ladder dewatering; the PIT tag detector was not activated during this interval.

Configuration

Configuration metadata includes details about the equipment used at the site to detect and record PIT tags. A new configuration sequence record is created when equipment is added or removed from the site and the period of time it was active. A configuration sequence without an End Date is currently active.

Select Configuration Number:
Start Date:  4/8/2006
End Date:  ----

Equipment History

Select a type to view the history of equipment installed at this site. A missing End Date implies the information about the equipment is current.

Select Equipment Type:
Equipment Code Description Start Date End Date

Photos


Event Log Summary

This displays the last 25 events, to see the full history, log in to the Dashboard and go to Interrogation Site > Event Logs.

Event Date Event Type Comments
6/2/2022 1:00:00 PM General

On site to replace D1's main air recirculating fan motor. The system had another flow failure alarm days before so the motor was deemed as the culprit. The air filters for the system were replaced for both units. A spare flow sensor, fan assembly with motor, and replacement filters were left for future use. D1 was running great when I left.

5/25/2022 1:00:00 PM General

On site to investigate a "No Flow" alarm for D1 on the dehumidification system. When I first started D1, the unit made a loud humming noise. I opened it up to find the main motor not blowing air but vibrating. After shaking the motor housing, it began to quiet down and begin blowing air. I stopped and started the unit many times and noticed the initial start of the unit was louder than the other unit but it did quiet down and begin blowing air as expected. The motor for the unit is to be replaced soon.

While on site, I replace D1's flow sensor. I tested the existing sensor and everything seemed to be fine with it. I replaced the sensor in case it was an intermittent problem with the sensor, I wanted to be able to eliminate that concern moving forward.

3/13/2022 8:50:00 AM SiteUp

Per site personnel the Channel was watered up on March 13th at 08:50

2/10/2022 9:44:00 AM General

It was noticed that files for M4 were not being submitted after the platform changeover last week. The automatic upload was disabled in the configuration. This was changed on both IPC PC's today.

2/3/2022 2:00:00 PM General

The antenna testing tool known as ARRT (automatic read range testing), was upgraded. The PLC and user interface to control the ARRT testing system was upgraded due to obsolete equipment with the older system. The new system also has some more features for future testing and is more user friendly for personnel running the test. 

11/29/2021 7:11:39 AM ErrorsCleared This check is being done once a day. It is part of the operation of the transceiver that cant be turned off.
11/27/2021 9:53:06 AM ErrorsCleared
11/26/2021 10:15:20 AM ErrorsCleared
11/24/2021 11:42:15 AM ErrorsCleared This check happened after power was restored on the system. It happens once a day.
11/8/2021 11:52:00 AM General
11/8/2021 10:58:00 AM General

Per Site Bio,

Yes, we have the B2CC locked out and tagged out for maintenance and repair. Part of the lockout process was to disable the PIT system, including the battery backup. I’m sorry this info was not relayed to you guys. I suspect this maintenance work will go for 2-3 weeks with absolutely no flow down the corner collector while people are in there with equipment

10/27/2021 7:40:56 AM ErrorsCleared Noted.
10/26/2021 7:43:31 AM ErrorsCleared Noted.
10/25/2021 8:04:48 AM ErrorsCleared Noted
10/24/2021 7:23:15 AM ErrorsCleared The system is checking for overcurrent once a day. Cant turn this feature off.
10/23/2021 8:47:09 AM ErrorsCleared Alarm generated during in water sealing and inspection
10/22/2021 7:07:25 AM ErrorsCleared This alarm will be investigated by PSMFC. The channel is dewatered at this time.
10/21/2021 7:34:11 AM ErrorsCleared System was locked out by the COE personnel to do some maintenance in the channel.
10/20/2021 7:12:32 AM ErrorsCleared Alarm was caused by site personnel doing maintenance in the channel.
10/19/2021 7:44:00 AM General

The site was locked out by the COE personnel this morning for maintenance work in the channel. Power should be restored this afternoon.

10/6/2021 7:23:39 AM ErrorsCleared Alarms were generated during water up and water down of the channel.
9/14/2021 8:39:46 AM ErrorsCleared Noted
9/1/2021 8:08:06 AM ErrorsCleared The channel was dewatered at 12:00 AM PST.
9/1/2021 12:00:00 AM SiteDown

The site was shut down for the season 9-1-2021 at 0000.

7/29/2021 2:00:00 PM General

Examination of secondary data shows that no detections occurred during the data gaps on 7-29-2021.

Archived Event Logs